Given the functions: f(x)=-7x g(x) = |x–51 h(x) = x+.5 Evaluate the function (4-3)(x) for x=2. (-8)(2) is
Answer #1

f(x) = -7x

g(x) = |x - 5|

Therefore,

(f - g)(x)

= f(x) - g(x)

= (-7x) - |x - 5|

Therefore,

(f - g)(x) = (-7x) - |x - 5|

Substituting x = 2 we get

(f - g)(2) = (-7(2)) - |2 - 5| = -14 - |-3| = -14 - 3 = -17

Therefore,

Answer: -17